OpenStudy (anonymous):

1. Solve: 3/4y−2=6−1/4y A. y = 4 B. y = 16 C. y = 2 D. y = 8

OpenStudy (anonymous):

3/4y -2 = 6 - 1/4y 3/4y + 1/4y = 6 + 2 solve that !!

OpenStudy (anonymous):

3/4y−2=6−1/4y 3/4y +1/4y = 6+2 4/4y = 8 y = 8